What is the increase in momentum if the kinetic energy is increased by 0.1% ?

Grade:11

1 Answers

Arun
6 years ago
 

Let k1 be the initial kinetic energy,k2 be the final kinetic energy.

 

let p1 be intial momentum & p2 be final momentum.

 

acc. to question,

 

k2=k1+0.1%k1

 

k2=k1+0.001k1=1.001k1

 

1\2(mV2)2=1.001(1\2mV12)

 

V2=(√1.001)V1

 

on mutiplying by mass m,

 

mV2=(√1.001)mV1

 

p2=(√1.001)p1.         now,

 

Δp=p2-p1             =>p1(√1.001-1)

 

%age of change in momentum=100(Δp\p1)

 

                                              =(1.00049-1)100

 

                                              =0.049%